Chinas inflation rate is 17.1% while Vietnams inflation rate is 2.6%. If the Chinese real exchange rate increased by 7.2%,
Answer #1

Change in real exchange rate = Change in Nominal Exchange rate * (Change in domestic inflation / Change in foreign Inflation)

(1 + 0.072) = Change in Nominal Exchange rate * [(1 + 0.171) / (1 + 0.026)]

Change in Nominal Exchange rate = (1.072 / 1.141) = 0.939 which means nominal exchange rate fell by 6.07%